Stock Issuance and Treasury Stock

Diva, Inc., recorded certain capital stock transactions shown in the following journal entries: (1) issued common stock for $20 cash per share, (2) purchased treasury shares at $25 per share, and (3) sold some of the treasury shares:

1 Cash 437,000
Common stock 43,700
Paid-in capital in excess of par value 393,300

2 Treasury stock 77,000
Cash 77,000

3 Cash 63,360
Treasury stock 52,800
Paid-in capital from treasury stock 10,560

a. How many shares were originally issued?

Answer

shares

b. What was the par value of the shares issued?

$Answer

per share

c. How many shares of treasury stock were acquired?

Answer

shares

d. How many shares of treasury stock were sold?

Answer

shares

e. At what price per share was the treasury stock sold?

$Answer

per share
Answer #1

Solution a:

Originally issued shares = Issued Proceeds / $20 per share = $437000 / 20 = 21,850 shares

Solution b:

Par value of the shares issued = Common Stock Value / Originally issued shares = $43700 / 21850 = $2 per share

Solution c:

Shares of treasury stock acquired = Treasury Stock Value acquired / $25 = $77000 / $25 = 3,080 shares

Solution d:

Shares of treasury stock sold = treasury stock value sold / $25 = $52,800 / $25 = 2,112 shares

Solution e:

Price per share, the treasury stock sold = Cash received on sale of treasury stock / Shares of treasury stock sold

= 63,360 / 2112 = $30 per share